Consider the situation of the previous problem. Find the average electric field energy stored in the capacitor and the average magnetic field energy stored in the coil.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol R = 300Ω C = 20 μF = 20 * 10^-6 F L = 1H, Z = 500 (from 14) Ε base 0 = 50 V, I base 0 = E base 0/Z = 50/500 = 0.1 A Electric Energy stored in Capacitor = (1/2) CV^2 = (1/2) × 20 × 10^–6 × 50 × 50 = 25 × 10^–3 J = 25 mJ Magnetic field energy stored in the coil = (1/2) L I base 0^2 = (1/2) * 1 * (0.1)^2 = 5 * 10^-3 J = 5 mJ